OpenStudy (anonymous):

LIMIT ( 2x - lxl )/ ( l3xl - 2x ) x-->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ 2x-\left| x \right| }{ \left| 3x \right|- 2x}\]

OpenStudy (anonymous):

to solve this you have to take side limits when x approaches 0 from negative and from positive if those limits are equal and is a real number then it exists hint: this limit doesnt exist

OpenStudy (anonymous):

uhmmm, there are absolute values there @ocelot40 . How do u go around them?

OpenStudy (anonymous):